OpenStudy (anonymous):

Hired for a job that lasts 31 days, payment plan is as follows; Plan: $0.01 for day 1 $0.02 for day 2 $0.04 for day 3 $0.08 for day 4 $0.16 for day 5 ….and so on every day is double the previous days salary, with no signing bonus. Create a function of the plan where t is the number of days, the function will represent what you earn each day. What is the function? Write it for 31 days.

OpenStudy (anonymous):

F(n) = 2^n -1 for n =number of days F(n) = total amount of money earned F(n) = 2^(n-1) for n= number of days F(n) = money earned that day

OpenStudy (anonymous):

I dont think that is right... day 3 says 0.08 but if I use that function 2^(3-1)=4, not what the question said it should be, 0.08...

OpenStudy (anonymous):

divide all answers by 100 :D sorry :D

OpenStudy (anonymous):

it's day 4 for 0.08 by the way :D

OpenStudy (anonymous):

Oh I see, that makes more sense I didnt read it right

OpenStudy (anonymous):

Thats still not right though. 2^(n-1) 2^5-1 = 2^4=16/100=2.58224 not the answer they provided

OpenStudy (anonymous):

5 days should be 0.16